# SemHiTok: A Unified Image Tokenizer via Semantic-Guided Hierarchical Codebook for Multimodal Understanding and Generation

### v1

@@ Abstract
We present SemHiTok, a unified image Tokenizer via Semantic-Guided Hierarchical codebook that provides consistent discrete feature representations for multimodal understanding and generation tasks. Recently, unified multimodal large models (MLLMs) for understanding and generation have sparked exploration within research community. Previous works attempt to train a unified image tokenizer by combining loss functions for semantic feature reconstruction and pixel reconstruction. However, due to the differing levels of features prioritized by multimodal understanding and generation tasks, joint training methods face significant challenges in achieving a good trade-off. SemHiTok addresses this challenge through Semantic-Guided Hierarchical codebook which builds texture sub-codebooks on pre-trained semantic codebook. This design decouples the training of semantic reconstruction and pixel reconstruction and equips the tokenizer with low-level texture feature extraction capability without degradation of high-level semantic feature extraction ability. Our experiments demonstrate that SemHiTok achieves state-of-the-art rFID score at $256\times 256$ resolution compared to other unified tokenizers, and exhibits competitive performance on multimodal understanding and generation tasks.
@@ 1 Introduction
To address this challenge, a straight forward solution is to train the semantic and texture vector quantizer separately, then concatenate quantized token either along the token length or across dimension. However, this approach increases the number of tokens thereby raising computational complexity, or results in a codebook that is too large to be suitable for MLLM. Inspired by the observation that image patches with the same semantic code tend to have similar texture feature, we propose SemHiTok, a unified image tokenizer that provides consistent feature representations for multimodal understanding and generation tasks through a unique hierarchical codebook design. The crucial point is to use same discrete features for multimodal understanding and generation tasks while decoupling the learning of semantic and texture features. Through semantic-guided hierarchical codebook, we achieve a unified representation of semantic and texture features for understanding and generation task. Unlike existing approaches that directly employ a single codebook to encode different level features, SemHiTok’s semantic-guided hierarchical codebook support stage-wise training paradigm where each stage exclusively optimizes specific hierarchy level features. This novel design allows us to achieve a better trade-off between semantic and texture feature extraction.
Our framework exhibits remarkable performance on multimodal understanding and generation tasks. For multimodal understanding, we surpass other next-token-based approaches using discrete visual inputs under the LLaVA-1.5 setting. SemHiTok also achieves competitive rFID metrics at a resolution of $256\times 256$ on ImageNet. For text-to-image generation tasks, we achieve state-of-the-art results, outperforming existing next-token autoregressive generation methods. These findings indicate that SemHiTok achieves a better trade-off between semantic and texture feature extraction capabilities, and demonstrates the great potential for application in MLLMs.
@@ 2.1 Image Tokenizer
Tokenization for Generation. Image tokenizers are crucial for autoregressive image generation[45, 38, 44]. VQVAE[45] learns a discrete representation using a learnable codebook in auto-encoder architectures. VQGAN[57] further improves better perceptual quality by using adversarial training[16]. advanced the existing architecture by integrating perceptual loss and discriminator loss, alongside adversarial loss, to enhance reconstruction quality. This approach yields more precise and detailed image representations, significantly improving upon previous methodologies in image generation and processing. Subsequently, ViT-VQGAN[58] and Efficient-VQGAN[5] advance the framework with the transformer design. In recent literature, researchers are turning to efficient codebook structures[39, 63, 58] and better quantization methods[61, 59] to improve generation performance and compression rates. IBQ[39] propose the Index Backpropagation Quantization codebook update method, achieving stable training of large-scale codebooks. TiTok[59] uses a self-attention mechanism to transfer image token information to latent tokens, achieving higher compression rates and improving generation performance. Although these methods efficiently retain low-level texture information, they frequently fail to capture high-level semantic information, which limits their application to multimodal understanding tasks.
@@ 3.3 Multimodal Downstream Tasks
Table 1: Comparison of reconstruction quality on the ImageNet-50k validation set. *: model is only trained on ImageNet. &: quantizer uses residual quantization (RQ), where the total Num are multiplied by RQ depth.
@@ 4.1 Experimental Setup
Datasets. The training of SemHiTok is conducted in two stages. During SPC training, we train the semantic tokenizer for one epoch on 50M subset of COYO-700M[4]. For PRE stage, we first trained the model on ImageNet[9], and then fine-tune on 50M COYO-700M subset to improve its generalization, following LlamaGen[40]. To further improve the reconstruction and generation performance, we enlarge the size of the pixel decoder and fine-tune it on the same data. For multimodel understanding, we train VLM on LLaVA-Pretrain558K and LLaVA-v1.5-mix-665K. To evaluate the performance of the VLM, we test the model on vision-language benchmarks such as SEEDB: SEED-Image[21], POPE[27], MMMU[60], MMB[32], MME-P: MME-Perception[12] and MME[12]. We run the evaluation based on lmms-eval[64]. For visual generation with SemHiTok, we use 15M in-house text-image pairs,and evaluate on GenEval[15] and MJHQ30K[23].
@@ 4.2 Unified Image Tokenizer
Model Text Pretrain Res. GenEval MJHQ30K
Overall $\uparrow$ gFID $\downarrow$
SD v1.5 [37] CLIP ViT-L/14 512 0.43
SD v2.1 [37] CLIP ViT-H/14 768 0.50
SDXL [33] CLIP ViT-bigG 1024 0.55 9.55
PixArt-alpha [6] Flan-T5-XXL 512 0.48 6.14
DALL-E 3 [3] Flan-T5-XXL 1024 0.67
Show-o [54] Phi-1.5 256 0.53 15.18
Transfusion [65] 256 0.63
LWM [31] 256 17.77
LlamaGen [40] Flan-T5-XL 512 0.32
EMU3 [48] 512 0.54
Janus [48] 384 0.61 10.10
VILA-U [52] 256 12.81
SemHiTok(ours) 256 0.66 11.0
Table 3: Comparison of generation quality on GenEval [15] and MJHQ30K [23].
@@ 4.4 Text-to-Image Generation
The results of SemHiTok on the text-to-image generation task are shown in Tab. 3. On GenEval, SemHiTok surpasses all autoregressive generation models and most diffusion models, such as SDXL, and is very close to the best-performing DALL-E. This indicates that our generation model is able to align with the prompt at the instance level. We attribute this to the powerful text feature extraction capabilities of the Qwen2.5 LLM, as well as the SGHC, which contains semantic information and can efficiently align the text feature space, ensuring accuracy in object generation. On the MJHQ30K, our model achieved 11.0 gFID and also demonstrated competitive performance, exceeds show-o and VILA-U. In Fig. 6, we present some generated images from MJHQ30K. Our model also demonstrates consistency with the text on MJHQ30K, even with long prompts. Owing to the powerful reconstruction capabilities of SemHiTok, the generated images are rich in detail and highly realistic, such as the fur of human and animal, the texture of leaves, and so on.

### v3

@@ Abstract
We present SemHiTok, a unified image Tokenizer via Semantic-Guided Hierarchical codebook that provides consistent discrete feature representations for multimodal understanding and generation tasks. Recently, unified multimodal large models (MLLMs) for understanding and generation have sparked exploration within research community. Previous works attempt to train a unified image tokenizer by combining loss functions for semantic feature reconstruction and pixel reconstruction. However, due to the differing levels of features prioritized by multimodal understanding and generation tasks, joint training methods face significant challenges in achieving a good trade-off. SemHiTok addresses this challenge through Semantic-Guided Hierarchical codebook which builds texture sub-codebooks on pre-trained semantic codebook. This design decouples the training of semantic reconstruction and pixel reconstruction and equips the tokenizer with low-level texture feature extraction capability without degradation of high-level semantic feature extraction ability. Our experiments demonstrate that SemHiTok achieves excellent rFID score at $256\times 256$ resolution compared to other unified tokenizers, and exhibits competitive performance on multimodal understanding and generation tasks.
@@ 1 Introduction
To address this challenge, a straight forward solution is to train the semantic and texture vector quantizer separately, then concatenate quantized token either along the token length or across dimension[3]. However, this approach increases the number of tokens thereby raising computational complexity, or results in a codebook that is too large to be suitable for MLLM. Inspired by the observation that image patches with the same semantic code tend to have similar texture feature, we propose SemHiTok, a unified image tokenizer that provides consistent feature representations for multimodal understanding and generation tasks through a unique hierarchical codebook design. The crucial point is to use same discrete features for multimodal understanding and generation tasks while decoupling the learning of semantic and texture features. Through semantic-guided hierarchical codebook, we achieve a unified representation of semantic and texture features for understanding and generation task. Unlike existing approaches that directly employ a single codebook to encode different level features, SemHiTok’s semantic-guided hierarchical codebook support stage-wise training paradigm where each stage exclusively optimizes specific hierarchy level features. This novel design allows us to achieve a better trade-off between semantic and texture feature extraction.
Our framework exhibits remarkable performance on multimodal understanding and generation tasks. For multimodal understanding, we surpass other next-token-based approaches using discrete visual inputs under the LLaVA-1.5 setting. SemHiTok also achieves competitive rFID metrics at a resolution of $256\times 256$ on ImageNet. For text-to-image generation tasks, we achieve competitive results, compared existing next-token autoregressive generation methods. These findings indicate that SemHiTok achieves a better trade-off between semantic and texture feature extraction capabilities, and demonstrates the great potential for application in MLLMs.
@@ 2.1 Image Tokenizer
Tokenization for Generation. Image tokenizers are crucial for autoregressive image generation[45, 38, 44]. VQVAE[45] learns a discrete representation using a learnable codebook in auto-encoder architectures. VQGAN[57] further improves better perceptual quality by using adversarial training[16]. advanced the existing architecture by integrating perceptual loss and discriminator loss, alongside adversarial loss, to enhance reconstruction quality. This approach yields more precise and detailed image representations, significantly improving upon previous methodologies in image generation and processing. Subsequently, ViT-VQGAN[58] and Efficient-VQGAN[5] advance the framework with the transformer design. In recent literature, researchers are turning to efficient codebook structures[39, 63, 58, 3] and better quantization methods[61, 59] to improve generation performance and compression rates. IBQ[39] propose the Index Backpropagation Quantization codebook update method, achieving stable training of large-scale codebooks. Although these methods efficiently retain low-level texture information, they frequently fail to capture high-level semantic information, which limits their application to multimodal understanding tasks. FQGAN[3] uses multiple codebooks and product quantization where each codebook encodes a different type of feature (Pixel, CLIP[36], DINO[6] etc.). However, this work focuses only on image reconstruction and generation tasks.
@@ 3.3 Multimodal Downstream Tasks
Table 1: Comparison of reconstruction quality on the ImageNet-50k validation set. *: model is only trained on ImageNet. &: quantizer uses residual quantization (RQ), where the total Code Num are multiplied by RQ depth. : quantizer uses multiple codebooks and product quantization.
@@ 4.1 Experimental Setup
Datasets. The training of SemHiTok is conducted in two stages. During SPC training, we train the semantic tokenizer for one epoch on 50M subset of COYO-700M[4]. For PRE stage, we first trained the model on ImageNet[10], and then fine-tune on 50M COYO-700M subset to improve its generalization, following LlamaGen[40]. To further improve the reconstruction and generation performance, we enlarge the size of the pixel decoder and fine-tune it on the same data. For multimodel understanding, we train VLM on LLaVA-Pretrain558K and LLaVA-v1.5-mix-665K. To evaluate the performance of the VLM, we test the model on vision-language benchmarks such as SEEDB: SEED-Image[21], POPE[28], MMMU[60], MMB[33], MME-P: MME-Perception[13] and MME[13]. We run the evaluation based on lmms-eval[64]. For visual generation with SemHiTok, we use 15M in-house text-image pairs,and evaluate on GenAI[23] and MJHQ30K[24].
@@ 4.2 Unified Image Tokenizer
Model Params Res. GenAI-Bench MJHQ30K
Basic $\uparrow$ Advanced $\uparrow$ gFID $\downarrow$
PixArt-$\alpha$ [7] 0.6B 1024 6.14
SDXL [34] 2.6B 1024 0.83 0.63 9.55
LWM [32] 7B 256 0.63 0.53 17.77
Show-o [54] 1.5B 256 0.70 0.60 15.18
VILA-U [52] 7B 256 0.76 0.64 12.81
VILA-U [52] 7B 384 0.73 0.61 7.69
ILLUME [52] 7B 512 0.75 0.60 7.76
Janus [50] 1.3B 384 9.55
SemHiTok(ours) 3B 256 0.74 0.58 9.4
Table 3: Comparison of generation quality on GenAI [23] and MJHQ30K [24].
@@ 4.4 Text-to-Image Generation
The results of SemHiTok on the text-to-image generation task are shown in Tab. 3. On GenAI, SemHiTok demonstrates competitive performance under basic and advanced prompts. This indicates that our generation model is able to be aligned to complex prompt words. We attribute this to the powerful text feature extraction capabilities of the Qwen2.5 LLM, as well as the SGHC, which contains semantic information and can efficiently align the text feature space. On the MJHQ30K, our model achieved 9.4 gFID and also demonstrated competitive performance, exceeds show-o and VILA-U under same resolution. In Fig. 6, we present some generated images from MJHQ30K. Our model also demonstrates consistency with the text on MJHQ30K, even with long prompts. Owing to the powerful reconstruction capabilities of SemHiTok, the generated images are rich in detail and highly realistic, such as the fur of human and animal, the texture of leaves, and so on.